# Geodetic Exploration of the Interior Structure of Enceladus

### v1

@@ Abstract
Enceladus is among the most intriguing bodies in the solar system due to its astrobiological potential. Determining the extent and duration of sustainable habitability requires characterizing the interior properties, and in particular, the extent and distribution of the tidal heating.The intensity of any inferred geophysical activity in the core has direct implications for understanding the hydrothermal activity and the budget of biologically useful species. We build a statistical framework using an MCMC approach to constrain the interior based on currently available observations of libration, shape, heat budget, and mass. We use this framework to examine the extent that geodetic measurements can constrain the interior structure, with an emphasis on determining the partitioning of tidal dissipation between the shell and the core. We quantify plausible ranges of gravitational and displacement tidal Love numbers consistent with existing observations. We demonstrate that measuring the amplitude and phase of $k_{2}$ alone can only constrain the total tidally dissipated energy, without a sensitivity to the distribution of heat. However, measuring the amplitude and phase of $h_{2}$ or $l_{2}$ facilitates determination of the shell rheology. We provide the precisions required for measuring $k_{2}$, $h_{2}$, and/or $l_{2}$ that enable distinguishing between the two main scenarios of tidal heating, i.e., in the shell versus the core. We also investigate the effect of the structural heterogeneities of the shell on the tidal response. Lastly, we evaluate the efficacy of a suite of future geodetic measurements to constrain key interior properties essential to instantaneous and sustained habitability at Enceladus.
@@ 1 Introduction and Background
Flybys of Saturn’s moon Enceladus by the Cassini spacecraft revealed multiple intriguing features suggesting an active surface and interior: plumes that continuously erupt from a region approximately centered on the satellite’s south pole (Porco et al., 2006) which feed Saturn’s diffuse E-ring with water-ice crystals (Kempf et al., 2010); a very bright surface (Hendrix et al., 2010); a high surface heat flux; and, importantly, evidence for a subsurface ocean (Porco et al., 2006; Spencer et al., 2006; Howett et al., 2011; McKinnon, 2015; Thomas et al., 2016; Park et al., 2024). This subsurface water body likely meets the criteria for instantaneous habitability (Cockell et al., 2024): An extensive liquid water reservoir containing salts and complex organic compounds and the expected chemical disequilibrium required for supporting metabolism. These conditions can occur through water-rock interactions, presumably in inferred hydrothermal systems at or below the seafloor (Postberg et al., 2009a; Howett et al., 2011; Glein et al., 2015; Hsu et al., 2015; Cable et al., 2021; Cockell et al., 2024). These lines of evidence for a potentially habitable environment position Enceladus as a compelling target for future exploration by orbiters and landers (NASEM, 2023). Studying to what extent these requirements have been sustained at Enceladus requires understanding the geophysical processes in the interior over geological timescales (see the discussions of sustained or dynamical habitability in NASEM (2023); Cockell et al. (2024)).
@@
Analysis of Cassini’s imaging, thermal mapping, and Deep Space Network (DSN) radiometric tracking data explored the global shape, gravity field, libration (i.e., diurnal wobble of the ice shell), and the surface topography of Enceladus, providing insights into the interior layers’ properties. The large amplitude of forced physical libration suggests that the ice crust is mechanically decoupled from the rocky core, likely by a global ocean (Nadezhdina et al., 2016; Thomas et al., 2016; Van Hoolst et al., 2016; Park et al., 2024). The presence of a global ocean is also supported by analysis of the low-order gravity field (McKinnon, 2015; Čadek et al., 2016). Estimates of the ice shell thickness and the surface temperatures require a total heat loss of approximately 25-40 GW (Hemingway & Mittal, 2019; Nimmo et al., 2023), while the measured heat output from the South Polar Terrain (SPT) alone is 4-19 GW (Howett et al., 2011). With an estimated radiogenic heating rate of less than 0.3 GW (Vance et al., 2007; Roberts & Nimmo, 2008) and minimal additional heating due to libration (Rambaux et al., 2010; Shao & Nimmo, 2022), the most likely heat source is therefore dissipation due to the eccentricity tides exerted by Saturn (e.g., Nimmo et al., 2023). Initial analyses suggested that given the global properties of Enceladus and Saturn, tidal dissipation is insufficient to explain the estimated heat loss (Meyer & Wisdom, 2007). However, later studies demonstrated that Saturn is supplying tidal energy and angular momentum to its satellites at a higher rate than previous estimates (Lainey et al., 2017), due to resonance locking between the moons and the planet’s internal oscillation modes (Fuller et al., 2016; Lainey et al., 2017, 2020; Jacobson, 2022).
@@
It remains unknown whether tidal dissipation primarily occurs in Enceladus’s core or ice shell. Distinguishing between the existing end-member scenarios is vital for assessing Enceladus’s sustained habitability (e.g., NASEM, 2023; Cockell et al., 2024). A tidally active core can facilitate hydrothermal activity at the bottom of the ocean and the continuous contact between water and fresh rock, potentially supporting the chemical reactions necessary for life to develop and evolve over geological timescales (Vance et al., 2007; Hsu et al., 2015; Choblet et al., 2017). Therefore, a comprehensive description of the global heat budget allows the distinction between a primordial and a present-day hydrothermal source of organics, volatiles, and nanosilica materials in the plumes (Vance et al., 2007; Hsu et al., 2015; Choblet et al., 2017). Studies have considered viscous dissipation in the ice shell as the mechanism of maintaining the global ocean (Roberts & Nimmo, 2008; Shoji et al., 2013; Běhounková et al., 2015; Efroimsky, 2018; Kang & Flierl, 2020). These models require a mean viscosity of $\eta_{shell}\approx 10^{13-15}$ Pa.s which is lower than the estimated ice viscosity at the temperatures in Enceladus. Alternative studies propose dissipation in water-filled fractures (Kite & Rubin, 2016), in the ocean (Tyler, 2020), or in a highly porous and dissipative rocky core with a very low effective viscosity (Roberts, 2015; Choblet et al., 2017; Neumann & Kruse, 2019; Souček et al., 2019; Liao et al., 2020; Aygün & Čadek, 2023).
@@
Considering the ambiguous understanding of the interior structure, geophysical investigations of Enceladus must account for the possibility of heating in the deep interior. Ermakov et al. (2021) and Genova et al. (2024) propose an effective approach to explore the interior by measuring the potential tidal Love number $k_{2}$ along with the libration or gravity-topography admittance to constrain the interior structure (i.e., thicknesses and densities of layers) at Enceladus. However, the ranges of interior parameters considered in previous studies inherently assume negligible tidal dissipation in the core without providing a quantitative approach to discriminating between the different scenarios of tidal dissipation. As we show later, a tidally active core will affect the time-dependent gravity field of the body and produce a gravitational signal which resembles that of a tidally dissipative ice shell, resulting in a non-unique interpretation of the amplitude and phase of $k_{2}$. Also, gravity and topography measurements have revealed structural heterogeneities in the Enceladus ice shell (Yin & Pappalardo, 2015; Park et al., 2024; Schenk & McKinnon, 2024). Deviations of the body from spherical symmetry can potentially affect its tidal response (e.g., Dmitrovskii et al., 2022; Berne et al., 2023; Goossens & Sabaka, 2025) by driving deformation at higher harmonic degrees (e.g., $l=3,4$), and mode coupling which is expected to produce variations in degree-2 Love numbers evaluated at different spherical harmonic orders, i.e., $k_{2m}$ and $h_{2m}$ (e.g., Bagheri et al., 2021; Pou et al., 2022; Berne et al., 2023).
@@
Here, we revisit the range of the interior model parameters based on the currently available observations using a MCMC statistical framework. We use this framework to examine the effectiveness of a suite of geodetic measurements including static gravity field, complex tidal Love numbers, libration, and heat loss rate that can constrain the interior of Enceladus. We highlight the degeneracy in interpreting $k_{2}$ to locate the tidally active region of Enceladus and provide a roadmap to distinguish between the two classes of tidal dissipation. We use our framework to compute the required precisions for measuring the tidal and displacement Love numbers. Using the proposed suite of future geophysical measurements and synthetic observations that can become available in future missions, we consider constraining interior parameters, such as the thickness, density, and rheology of the interior layers. These constraints can significantly improve our understanding of the instantaneous and sustained habitability at Enceladus.
@@ 2 Methods
We describe the geophysical models needed to compute the geodetic quantities in our forward model. The models include the tidal response in the form of the complex Love numbers and maximum surface deformations, the viscoelastic models used in computing the tidal response and physical libration, and calculation of the static gravity coefficients. Detailed derivations are provided in the Appendix.
@@ 2.1 Tidal Response
The eccentric orbit of Enceladus around Saturn drives periodic gravitational tidal forcing of the moon. This forcing potential, truncated at spherical harmonic degree 2 and to first order for eccentricity, can be written as (Murray & Dermott, 1999):
@@
where $R_{E}$ is the radius of Enceladus, and $\Delta R$, $\Delta E$, and $\Delta N$ are the radial, easterly, and northerly displacements on the surface, respectively. Because of the viscoelastic behavior of materials, the tidal Love numbers are complex values, implying that there is a phase lag between the forcing potential and the tidal response. We compute the total heat generated by tides from the complex $k_{2}$ (Segatz et al., 1988):
@@
Here, $\epsilon_{k_{2}}$ is the phase lag of the perturbation in the gravitational potential, which can represent the total tidal energy generated throughout the body. In most cases, the tidal quality factor $Q$ is simply defined as $\tan(\epsilon_{k_{2}})$. However, this simplification is only correct when the phase lag is small such that $\tan(\epsilon_{k_{2}})\approx\sin(\epsilon_{k_{2}})$. For bodies like Enceladus and other satellites that may experience intense tidal heating and hence high tidal phase lag, this approximation is no longer valid (Efroimsky, 2012a, b). Here, we explicitly consider phase lags, as opposed to the quality factor $Q$. This choice is appropriate since for Enceladus the phase lags are potentially large, $Q$ is defined inconsistently in the literature, and our focus is on direct observables. We also generally assume that the total heat loss is equal to the tidally generated heat computed by Equation (5); in other words, we assume that Enceladus is currently in a thermal steady state.
@@ 2.2 Viscoelastic model
In order to predict the tidal response of Enceladus, we must consider rheological models that incorporate shear dissipation in rock and icy regions. When both viscous and elastic mechanisms are active, the viscoelastic response depends on the shear modulus $\mu$, frequency of excitation $\omega$, and the effective viscosity $\eta$. The Maxwell viscoelastic model provides a reasonable description of a material’s behavior at time scales that are significantly shorter or longer than the Maxwell (relaxation) time of the body defined as $\tau_{M}=\eta/\mu$. $\tau_{M}$ represents the approximate timescale at which the behavior transitions from being dominated by the elastic regime to the viscous regime. However, a Maxwell model may not be appropriate for describing tidal dissipation in planetary bodies that have tidal periods ranging from hours to a few days, which may be near the relaxation time of the body. Thus, applicability of Maxwell’s model for capturing tidal dissipation has been questioned for multiple planetary bodies (Bills et al., 2005; Castillo-Rogez et al., 2011; Steinbrügge et al., 2018; Bagheri et al., 2019; Petricca et al., 2024). Several models based on micro-scale physical processes and laboratory measurements have been proposed to mimic the behavior of a material at time scales between the elastic and viscous regimes (e.g., Castillo-Rogez et al., 2011; Bagheri et al., 2019, 2021; Petricca et al., 2024; Tobie et al., 2025). In particular, the Andrade model has been recommended to model tidal dissipation in celestial bodies taking into account the effect of anelastic behavior in the transition from dominantly elastic to dominantly viscous regimes (Castillo-Rogez et al., 2011; Renaud & Henning, 2018; Padovan et al., 2014; Bagheri et al., 2019, 2022a). A detailed discussion of various viscoelastic models can be found in Jackson & Faul (2010); Faul & Jackson (2015); Bagheri et al. (2019). We consider both Maxwell and Andrade viscoelastic models because of the potential proximity of the relaxation time of the ice shell to Enceladus’s orbital period, but most of our results assume the Andrade rheology.

### v3

# Exploring the Interior Structure and Mode of Tidal Heating in Enceladus

@@ Abstract
Enceladus is among the most intriguing bodies in the solar system due to its astrobiological potential. Determining the extent and duration of habitability (i.e., sustained habitability) at Enceladus requires characterizing its interior properties as well as the level and distribution of tidal heating. Inferring the intensity of geophysical activity in the core has direct implications for understanding the potential for hydrothermal activity and the supply of chemical species important for habitability to the ocean. We build a statistical framework using an MCMC approach to constrain the interior structure based on currently available estimates of libration, shape, heat flux, and total mass. We use this framework to examine the extent to which geodetic measurements can improve our understanding of the interior structure, with an emphasis on determining the partitioning of tidal dissipation between the shell and the core. We quantify plausible ranges of gravitational ($k_{2}$) and displacement ($h_{2}$ and $l_{2}$) tidal Love numbers consistent with existing observations. We demonstrate that measuring the amplitude and phase of $k_{2}$ alone can only constrain the total tidally dissipated energy, but not its radial distribution. However, measuring the amplitude and phase of $h_{2}$ or $l_{2}$ facilitates determining the extent of tidal dissipation in the shell and the core. We provide the precisions required for measuring $k_{2}$, $h_{2}$, and/or $l_{2}$ that enable distinguishing between the two main scenarios of tidal heating, i.e., in the shell versus the core. We also briefly explore the effect of the structural heterogeneities of the shell on the tidal response. Lastly, we evaluate the efficacy of a suite of future geodetic measurements to constrain key interior properties essential to understand the present-day (i.e., instantaneous) and long-term (sustained) habitability at Enceladus.
@@ 1 Introduction and Background
Flybys of Saturn’s moon Enceladus by the Cassini spacecraft revealed multiple intriguing features suggesting an active surface and interior: plumes that continuously erupt from a region approximately centered on the satellite’s south pole (Porco et al., 2006) which feed Saturn’s diffuse E-ring with water-ice crystals (Kempf et al., 2010); a very bright surface (Hendrix et al., 2010); a high surface heat flux (Spencer et al., 2006; Howett et al., 2011); and, importantly, evidence for a subsurface ocean (Porco et al., 2006; Iess et al., 2014; Thomas et al., 2016; Park et al., 2024). This subsurface water body likely meets the criteria for instantaneous habitability (Cockell et al., 2024): An extensive liquid water reservoir containing salts and complex organic compounds and the expected chemical disequilibrium required for supporting metabolism. These conditions can occur through water-rock interactions, presumably in inferred hydrothermal systems at or below the seafloor (Postberg et al., 2009; Howett et al., 2011; Hsu et al., 2015; Cable et al., 2021; Cockell et al., 2024). These lines of evidence for a potentially habitable environment position Enceladus as a compelling target for future exploration by orbiters and landers (NASEM, 2023). Investigating the extent to which these conditions might have been sustained at Enceladus requires understanding the geophysical processes in the interior over geological timescales (see Cockell et al. (2024) and NASEM (2023) for discussions of sustained or dynamical habitability).
@@
Analysis of Cassini’s imaging, thermal mapping, and radiometric tracking data explored the global shape, gravity field, libration (i.e., diurnal wobble of the ice shell), and the surface topography of Enceladus, providing insights into the interior layers’ properties. The large amplitude of forced physical libration suggests that the ice crust is mechanically decoupled from the rocky core, likely by a global ocean (Nadezhdina et al., 2016; Thomas et al., 2016; Van Hoolst et al., 2016; Park et al., 2024). The presence of a global ocean is also supported by analysis of the low-order gravity field (McKinnon, 2015; Čadek et al., 2016).
The total rate of heat loss at Enceladus remains poorly constrained (Howett et al., 2025). Estimates of the ice shell thickness and surface temperatures suggest a total heat loss of approximately 25-40 GW (Hemingway & Mittal, 2019; Nimmo et al., 2023), while the measured heat output from the South Polar Terrain (SPT) alone is estimated to be 4-19 GW (Howett et al., 2011). With an estimated radiogenic heating rate of less than 0.3 GW (Vance et al., 2007; Roberts & Nimmo, 2008; Nimmo et al., 2018) and minimal additional heating due to libration (Shao & Nimmo, 2022), the most likely heat source is therefore dissipation due to the eccentricity tides exerted by Saturn (e.g., Nimmo et al., 2023). Initial analyses suggested that given the global properties of Enceladus and Saturn, tidal dissipation is insufficient to explain the estimated heat loss (Meyer & Wisdom, 2007). However, later studies suggested that Saturn is supplying tidal energy and angular momentum to its satellites at a higher rate than previous estimates (Lainey et al., 2017), suggesting a resonance locking between the moons and the planet’s internal oscillation modes (Fuller et al., 2016; Lainey et al., 2017; Nimmo et al., 2018; Lainey et al., 2020; Jacobson, 2022). Alternatively, there may exist transient periods of tidal heating. In this scenario, the present-day ocean would be a manifestation of a cycle of ocean formation and freezing as a result of a dynamically unstable orbit (Meyer & Wisdom, 2008a; Goldreich et al., 2025), or instead as a result of changes in the mean motion resonance in the case of an ongoing feedback between the moon’s thermal and orbital evolution. (Ojakangas & Stevenson, 1986; Meyer & Wisdom, 2008b; Renaud et al., 2021; Goldreich et al., 2025). A transient ocean could also be the result of recent catastrophic events that affect the Saturnian system involving Enceladus (e.g., Asphaug & Reufer, 2013; Wisdom et al., 2022; Teodoro et al., 2023).
@@
It remains unknown whether tidal dissipation occurs mainly in Enceladus’s core or ice shell. Distinguishing between the existing end-member scenarios is vital for assessing Enceladus’s sustained habitability (e.g., NASEM, 2023; Cockell et al., 2024). A tidally active core can drive hydrothermal activity at the bottom of the ocean and the continuous contact between water and fresh rock, potentially supporting the chemical reactions necessary for life to develop and evolve over geological timescales (Vance et al., 2007; Hsu et al., 2015; Choblet et al., 2017). A comprehensive description of the global heat budget would make it possible to distinguish between a primordial and a present-day hydrothermal source of organics, volatiles, and nanosilica materials in the plumes. Studies have considered viscous dissipation in the ice shell as the mechanism to maintain the global ocean (e.g. Roberts & Nimmo, 2008; Shoji et al., 2013; Efroimsky, 2018; Kang & Flierl, 2020). These models require a mean viscosity of $\eta_{shell}\approx 10^{13-15}$ Pa.s which is significantly lower than the estimated ice viscosity at the temperatures expected in the outer shell of Enceladus (e.g. Goldsby & Kohlstedt, 2001). Alternative studies propose dissipation in water-filled fractures (Kite & Rubin, 2016) or in a highly porous and dissipative rocky core with a very low effective viscosity (Roberts, 2015; Choblet et al., 2017; Souček et al., 2019; Liao et al., 2020; Rovira-Navarro et al., 2022; Aygün & Čadek, 2023). Dissipation due to resonance effects in the ocean has also been proposed (Tyler, 2020), but was found incompatible with the estimated shell and ocean thicknesses (Hay & Matsuyama, 2017; Matsuyama et al., 2018).
@@
Considering the ambiguities in understanding the interior, geophysical investigations of Enceladus must account for the possibility of heating in the deep interior. Ermakov et al. (2021) and Genova et al. (2024) explore models of the interior by measuring the potential tidal Love number $k_{2}$ along with the libration or gravity-topography admittance to constrain the interior structure (i.e., thicknesses and densities of layers) at Enceladus. However, the ranges of interior parameters considered in these studies assume a nearly rigid core and thus inherently produce negligible tidal dissipation in the core. Therefore, these studies do not address the challenge of discriminating between the different scenarios of tidal dissipation. As we will show, a tidally active core will affect the time-dependent gravity field of the body and produce a gravitational signal that resembles that of a tidally dissipative ice shell, resulting in a non-unique interpretation of the amplitude and phase of $k_{2}$. In addition, gravity and topography measurements have revealed structural heterogeneities in the Enceladus ice shell (Yin & Pappalardo, 2015; Park et al., 2024; Schenk & McKinnon, 2024). Deviations of the body from spherical symmetry can potentially affect its tidal response (e.g., Zhong et al., 2012; Qin et al., 2014; Wahr et al., 2014; Lau et al., 2015; Dmitrovskii et al., 2022; Berne et al., 2023a) by driving deformation at higher harmonic degrees (e.g., $l=3,4$) through mode coupling which is expected to produce variations in degree-2 Love numbers evaluated at different spherical harmonic orders, i.e., $k_{2m}$ and $h_{2m}$ (e.g., Bagheri et al., 2021; Pou et al., 2022; Berne et al., 2023a).
Here, we review the physical aspects of the libration and tidal responses of Enceladus as a generic icy satellite. We explore the range of the interior model parameters compatible with the currently available observations using an MCMC statistical framework. We use this framework to examine the effectiveness of a suite of geodetic measurements including static gravity field, complex tidal Love numbers, libration, and heat loss rate that can constrain the interior of Enceladus.
@@
We highlight the degeneracy in interpreting $k_{2}$ to identify the tidally active region(s) of Enceladus and provide a roadmap to distinguish between the two classes of tidal dissipation. We use our framework to compute the required precisions for measuring the tidal and displacement Love numbers. This framework allows for an improved understanding of the sensitivity of the observables to each of the interior parameters. Using the proposed suite of future geophysical measurements and synthetic observations that could be available in future missions, we consider constraining interior parameters, such as the thickness, density, and rheology of the interior layers. These constraints can significantly improve our understanding of the parameters that control the instantaneous and sustained habitability at Enceladus. The methodology and derivations developed here apply broadly to other icy bodies.
@@ 2 Methods
We describe the geophysical models needed to compute the geodetic quantities in our forward model. The models include the tidal response in the form of the complex Love numbers and maximum surface deformations, the viscoelastic models used in computing the tidal response and physical libration, and calculation of the static gravity coefficients. Detailed derivations are provided in the appendices.
@@ 2.1 Tidal Response
The eccentric orbit of Enceladus around Saturn drives periodic gravitational tidal forcing of the moon. This forcing potential, truncated at spherical harmonic degree 2 and to first order in eccentricity, can be written as (Murray & Dermott, 1999):
@@
where $R_{E}$ is the radius of Enceladus, and $\Delta R$, $\Delta E$, and $\Delta N$ are the radial, easterly, and northerly displacements on the surface, respectively.
In general, for a body with lateral variations in structure, the spherical harmonic expansion of the tidal response depends on both the degree (here $l=2$) and the order $m$. Thus, whereas for a spherically symmetric body, $k_{22}=k_{20}=k_{2}$ and $h_{22}=h_{20}=h_{2}$, for a non-spherically symmetric body these quantities can be different. In this case, $k_{2m}=W_{2m}(r,\theta,\phi,t)/U_{2m}(r,\theta,\phi,t)$, and $h_{2m}=gH_{2m}(r,\theta,\phi,t)/U_{2m}(r,\theta,\phi,t)$, where $W_{2m}$ and $H_{2m}$ are the spherical harmonic expansion terms of the potential and surface displacement (assuming l=2 forcing only).
Because of the viscoelastic behavior of materials, the tidal Love numbers are complex values, implying that there is a phase lag between the forcing potential and the tidal response. The total heat generated by tides can be directly computed from the complex $k_{2}$ (Peale & Cassen, 1978; Segatz et al., 1988):
@@
Here, $\epsilon_{k_{2}}$ is the phase lag of the perturbation in the gravitational potential, which can represent the total tidal energy generated throughout the body. In most cases, the tidal quality factor $Q$ is simply defined as $\tan(\epsilon_{k_{2}})$. However, this simplification is only correct when the phase lag is small such that $\tan(\epsilon_{k_{2}})\approx\sin(\epsilon_{k_{2}})$. For bodies like Enceladus that may experience intense tidal heating and high tidal phase lag, this approximation is no longer valid (Efroimsky, 2012). Here, we explicitly consider phase lags, as opposed to the quality factor $Q$.
@@ 2.2 Viscoelastic Model
To predict the tidal response of Enceladus, we must consider rheological models that incorporate shear dissipation in rock and icy regions. When both viscous and elastic mechanisms are active, the viscoelastic response depends on the shear modulus $\mu$, frequency of excitation $\omega$, and the effective viscosity $\eta$. The Maxwell viscoelastic model provides a reasonable description of a material’s behavior at time scales that are significantly shorter or longer than the Maxwell (relaxation) time of the body defined as $\tau_{M}=\eta/\mu$. $\tau_{M}$ represents the approximate timescale at which the behavior transitions from being dominated by the elastic regime to the viscous regime. However, a Maxwell model may not be appropriate for describing tidal dissipation in planetary bodies that have tidal periods ranging from hours to a few days, which may be near the relaxation time of the body. Thus, applicability of Maxwell’s model for capturing tidal dissipation has been questioned for multiple planetary bodies (Bills et al., 2005; Castillo-Rogez et al., 2011; Bagheri et al., 2019; Petricca et al., 2024). Several models based on micro-scale physical processes and laboratory measurements have been proposed to mimic the behavior of a material in time scales between the elastic and viscous regimes (e.g., Castillo-Rogez et al., 2011; Bagheri et al., 2019, 2021; Petricca et al., 2024; Tobie et al., 2025). In particular, the Andrade model has been favored to model tidal dissipation in celestial bodies taking into account the effect of anelastic behavior in the transition from dominantly elastic to dominantly viscous regimes (Castillo-Rogez et al., 2011; Renaud & Henning, 2018; Padovan et al., 2014; Bagheri et al., 2019, 2022a). A detailed discussion of various viscoelastic models can be found in Jackson & Faul (2010); Faul & Jackson (2015); Bagheri et al. (2019). We consider both Maxwell and Andrade viscoelastic models because of the potential proximity of the relaxation time of the ice shell to Enceladus’s orbital period, but most of our results assume the Andrade rheology.